In an electric heater the inlet temperature is 15°C, water is flowing at the rate of 300 grams per minute. The voltmeter measuring voltage across the heating element reads 120 volts and an ammeter current taken reads 10 amperes. When steady state is finally reached, what is the final reading of the outlet thermometer?


1
Expert's answer
2021-10-07T10:40:02-0400

"Q=E"

"Q=cm(t_2-t_1)"

"E=UIt"


"cm_0t(t_2-t_1)=UIt\to t_2=UI/(cm_0)+t_1="


"=120\cdot10/(4200\cdot0.3)+15=15.95°" . Answer